# Parameter-Efficient Fine-Tuning of State Space Models

### v1

@@ C.3 Limitations of Applying Prompt-based Methods on SSMs
We provide the formal version of Proposition 1 and its corresponding proof here. We start by introducing the necessary notations. Denote the space of S4 mechanisms with $D$ channels as ${\mathcal{F}}_{\text{S4},D}$. Let ${\bm{H}}_{0}=({\bm{h}}_{0}^{(1)},{\bm{h}}_{0}^{(2)},\dots,{\bm{h}}_{0}^{(D)})\in{\mathbb{R}}^{H\times D}$ represent the initial hidden state, and ${\bm{X}}=({\bm{x}}_{1},{\bm{x}}_{2},\ldots,{\bm{x}}_{N})\in{\mathbb{R}}^{D\times N}$ denote the input sequence. The output of the S4 mechanism is represented as $f({\bm{X}};{\bm{H}}_{0})$. Furthermore, for $d$-th channel, let state transition matrix $\overline{{\bm{A}}}^{(d)}=\operatorname{diag}{(a_{1}^{(d)},\cdots,a_{H}^{(d)})}$ and input transition vector $\overline{{\bm{B}}}^{(d)}=(b_{1},\cdots,b_{H})^{\top}$, where $d=1,\ldots,D$. For any vector ${\bm{v}}\in{\mathbb{R}}^{n}$, we use ${\bm{v}}_{i:j}\in{\mathbb{R}}^{j-i}$ to denote the subvector of ${\bm{v}}$ containing elements from $i\in{\mathbb{N}}^{+}$ to $j\in{\mathbb{N}}^{+}$, where $i<j$. Similarly, for any matrix ${\bm{M}}\in{\mathbb{R}}^{m\times n}$, we use ${\bm{M}}_{i_{1}:j_{1},i_{2}:j_{2}}$ to denote the submatrix containing rows $i_{1}\in{\mathbb{N}}^{+}$ to $j_{1}\in{\mathbb{N}}^{+}$ and columns $i_{2}\in{\mathbb{N}}^{+}$ to $j_{2}\in{\mathbb{N}}^{+}$, where $i_{1}<j_{1}$, $i_{2}<j_{2}$.
Proposition 5 (Formal Version of Proposition 1).
@@ C.4 Optimal Application of LoRA in SSM-based Models
Several studies (Hu et al., 2023; He et al., 2021) present findings on Transformers, indicating that applying LoRA to linear projection matrices yields performance comparable to or marginally superior to that of attention layers. In contrast, our experimental results on SSMs reveal that applying LoRA to linear projection matrices is more effective than applying it to S6 (see Table 3). To elucidate this phenomenon, we examine the influence of updating linear projection matrices on the model’s output.
@@
For the feasibility of the analysis, we consider a simplified SSM-based architecture which only consists of the input projection matrix $\bm{W}_{\text{in}}\in{\mathbb{R}}^{D\times D}$ and the S6 module parameterized by diagonal state transition matrices $\{{\bm{A}}^{(d)}\}_{d=1}^{D}$ with ${\bm{A}}^{(d)}\in{\mathbb{R}}^{H\times H}$, the weight matrices $\bm{W}_{{\bm{B}}},\bm{W}_{{\bm{C}}}\in{\mathbb{R}}^{H\times D}$ for computing input-dependent input transition vectors ${\bm{B}}_{n}\in{\mathbb{R}}^{H}$ and output mapping vectors ${\bm{C}}_{n}\in{\mathbb{R}}^{H}$, the down and up projection matrices ${\bm{W}}_{\bm{\Delta},\downarrow}\in{\mathbb{R}}^{D\times R},{\bm{W}}_{\bm{\Delta},\uparrow}\in{\mathbb{R}}^{R\times D}$ (where $R$ is the rank) for low-rank weight matrices for computing the input-depdenent step size $\bm{\Delta}_{n}=(\Delta^{(1)}_{n},\ldots,\Delta^{(D)}_{n})\in{\mathbb{R}}_{D}$, for $n=1,\ldots,N$. Define $\bm{W}_{\text{S6}}=[\bm{W}_{{\bm{B}}}^{\top},\bm{W}_{{\bm{C}}}^{\top},{\bm{W}}_{\bm{\Delta},\uparrow}^{\top}]^{\top}\in{\mathbb{R}}^{(2H+R)\times D}$. In the Mamba implementation, $\bm{W}_{\text{S6}}$ is implemented as the weight matrix of a single linear layer, referred to as x_proj in the codebase. Therefore, the parameters of the S6 can be formulated as
${\bm{\theta}}(\cdot;\{{\bm{A}}\}_{d=1}^{D},\bm{W}_{\text{S6}},{\bm{W}}_{\bm{\Delta},\downarrow},\bm{W}_{\text{in}})=\{\overline{{\bm{A}}}_{n},\overline{{\bm{B}}}_{n},{\bm{C}}_{n}\}_{n=1}^{N}.$ (20)
Consider input sequence ${\bm{X}}=({\bm{x}}_{1},\ldots,{\bm{x}}_{N})\in{\mathbb{R}}^{D\times N}$. Let ${\bm{Z}}=({\bm{z}}_{1},\ldots,{\bm{z}}_{N})\in{\mathbb{R}}^{D\times N}$ denote the intermediate output after the input projection. The intermediate output at position $n\in\{1,\ldots,N\}$ is
${\bm{z}}_{n}=\bm{W}_{\text{in}}\bm{x}_{n}.$ (21)
Note that
${\bm{B}}_{n}=\bm{W}_{{\bm{B}}}{\bm{z}}_{n},\quad{\bm{C}}_{n}=\bm{W}_{{\bm{C}}}{\bm{z}}_{n},\quad\bm{\Delta}_{n}=\operatorname{softplus}({\bm{W}}_{\bm{\Delta},\uparrow}{\bm{W}}_{\bm{\Delta},\downarrow}{\bm{z}}_{n}+{\bm{\beta}}_{\bm{\Delta}}),$ (22)
and after discretization, we have
$\overline{{\bm{A}}}_{n}^{(d)}=\exp(\Delta_{n}^{(d)}{\bm{A}}^{(d)}),\quad\overline{{\bm{B}}}_{n}=\Delta_{n}^{(d)}{\bm{B}}_{n}=\Delta_{n}^{(d)}\bm{W}_{{\bm{B}}}{\bm{z}}_{n}.$ (23)
Combining (21), (22) and (23) yields
${\bm{\theta}}({\bm{X}};\{{\bm{A}}\}_{d=1}^{D},\bm{W}_{\text{S6}},{\bm{W}}_{\bm{\Delta},\downarrow},\bm{W}_{\text{in}})=\{\overline{{\bm{A}}}_{n},\overline{{\bm{B}}}_{n},{\bm{C}}_{n}\}_{n=1}^{N},\text{ where}$ (24)
$\overline{{\bm{A}}}_{n}^{(d)}=\exp(\Delta_{n}^{(d)}{\bm{A}}^{(d)}),\quad\overline{{\bm{B}}}_{n}=\Delta_{n}^{(d)}\bm{W}_{{\bm{B}}}\bm{W}_{\text{in}}\bm{x}_{n},\quad{\bm{C}}_{n}=\bm{W}_{{\bm{C}}}\bm{W}_{\text{in}}\bm{x}_{n},$ (25)
$\bm{\Delta}_{n}=\operatorname{softplus}({\bm{W}}_{\bm{\Delta},\downarrow}{\bm{W}}_{\bm{\Delta},\uparrow}\bm{W}_{\text{in}}\bm{x}_{n}+{\bm{\beta}}_{\bm{\Delta}}).$ (26)
@@
In the following theorem, we demonstrate that applying LoRA exclusively to $\bm{W}_{\text{in}}$ is equivalent to applying it to $\bm{W}_{\text{S6}}$.
Lemma 6 (Detailed Version of Lemma 2).
Consider a model consists of an S6 module augmented with a linear input projection $\bm{W}_{\text{in}}\in{\mathbb{R}}^{D\times D}$. For any fine-tuned model where only $\bm{W}_{\text{S6}}$ is updated to $\overline{\bm{W}}_{\text{S6}}$, there exists $\widehat{{\bm{W}}}_{\text{in}}$ such that updating only $\bm{W}_{\text{in}}$ to $\widehat{{\bm{W}}}_{\text{in}}$ yields:
${\bm{\theta}}({\bm{X}};\{{\bm{A}}^{(d)}\}_{d=1}^{D},\overline{\bm{W}}_{\text{S6}},{\bm{W}}_{\bm{\Delta},\downarrow},\bm{W}_{\text{in}})={\bm{\theta}}({\bm{X}};\{{\bm{A}}^{(d)}\}_{d=1}^{D},\bm{W}_{\text{S6}},{\bm{W}}_{\bm{\Delta},\downarrow},\widehat{{\bm{W}}}_{\text{in}})$ (27)
@@
In this proof, we use $\overline{\cdot}$ to denote the corresponding notations for the model with only $\bm{W}_{\text{S6}}$ updated, and use $\widehat{\cdot}$ to denote the corresponding notations for the model with only $\bm{W}_{\text{in}}$ updated.
To demonstrate (27), it is sufficient, according to (24), to find $\widehat{{\bm{W}}}_{\text{in}}$ that satisfies the following equations:
$\overline{{\bm{W}}}_{{\bm{C}}}\bm{W}_{\text{in}}=\bm{W}_{{\bm{C}}}\widehat{{\bm{W}}}_{\text{in}}$ (28)
$\overline{{\bm{W}}}_{\bm{\Delta},\uparrow}\bm{W}_{\text{in}}={\bm{W}}_{\bm{\Delta},\uparrow}\widehat{{\bm{W}}}_{\text{in}}$ (29)
$\overline{{\bm{W}}}_{{\bm{B}}}\bm{W}_{\text{in}}=\bm{W}_{{\bm{B}}}\widehat{{\bm{W}}}_{\text{in}}.$ (30)
Since $\bm{W}_{\text{S6}}=\begin{bmatrix}\bm{W}_{{\bm{B}}}\\ \bm{W}_{{\bm{C}}}\\ {\bm{W}}_{\bm{\Delta},\uparrow}\end{bmatrix}$, the three conditions (28) can be written as
$\overline{\bm{W}}_{\text{S6}}\bm{W}_{\text{in}}=\bm{W}_{\text{S6}}\widehat{{\bm{W}}}_{\text{in}}.$ (31)
By applying Singular Value Decomposition (SVD) to $\bm{W}_{\text{S6}}$ and $\left(\bm{W}_{\text{S6}}-\overline{\bm{W}}_{\text{S6}}\right)\bm{W}_{\text{in}}$, we obtain:
@@
$\left(\bm{W}_{\text{S6}}-\overline{\bm{W}}_{\text{S6}}\right)\bm{W}_{\text{in}}={\bm{U}}^{\prime}\begin{bmatrix}{\bm{\Sigma}}^{\prime}&{\bm{O}}_{(2H+R)\times(D-2H-R)}\end{bmatrix}{\bm{V}}^{\prime\top},$ (33)
where ${\bm{U}},{\bm{U}}^{\prime}\in{\mathbb{R}}^{(2H+R)\times(2H+R)}$, ${\bm{\Sigma}},{\bm{\Sigma}}^{\prime}\in{\mathbb{R}}^{(2H+R)\times(2H+R)}$, and ${\bm{V}},{\bm{V}}^{\prime}\in{\mathbb{R}}^{D\times D}$. The diagonal elements of ${\bm{\Sigma}}$ and ${\bm{\Sigma}}^{\prime}$ are in decreasing order.
We let
$\bm{\widehat{W}}_{\textrm{in}}={\bm{V}}\begin{bmatrix}{\bm{\Sigma}}^{-1}{\bm{U}}^{\top}\overline{\bm{W}}_{\text{S6}}\bm{W}_{\text{in}}\\ {\bm{Q}}\end{bmatrix},$ (34)
where ${\bm{Q}}\in{\mathbb{R}}^{(D-2H-R)\times D}$ is an arbitrary matrix to be determined later. Plugging (32) and(34) back to $\bm{W}_{\text{S6}}\widehat{{\bm{W}}}_{\text{in}}$ and simplifying results in
$\bm{W}_{\text{S6}}\widehat{{\bm{W}}}_{\text{in}}$ (35)
$={\bm{U}}\begin{bmatrix}{\bm{\Sigma}}&{\bm{O}}_{(2H+R)\times(D-2H-R)}\end{bmatrix}{\bm{V}}^{\top}{\bm{V}}\begin{bmatrix}{\bm{\Sigma}}^{-1}{\bm{U}}^{\top}\overline{\bm{W}}_{\text{S6}}\bm{W}_{\text{in}}\\ {\bm{Q}}\end{bmatrix}$ ((32) & (34)) (36)
$=\overline{\bm{W}}_{\text{S6}}\bm{W}_{\text{in}},$ (Simplifying) (37)
@@
Appendix D Details of Sec. 5: Selective Dimension Tuning
@@
Let ${\mathcal{P}}^{H}$ denote the set of all $H\times H$ permutation matrices. Given this formulation, we present our first analysis of parameter efficiency for the S4 model in the following lemma. This analysis is based on the parameters after necessary discretization $(\overline{{\bm{A}}},\overline{{\bm{B}}},{\bm{C}})$.
Lemma 0 (Essential Discretized Parameter Set for S4).
Consider the parameters after discretization, i.e., $\overline{{\bm{A}}},\overline{{\bm{B}}},{\bm{C}}$. To achieve functional equivalence between the updated model and the target model, i.e., $\hat{f}\equiv f^{\star}$, it is sufficient to tune the following number of parameters:
$\min_{{\bm{P}}\in{\mathcal{P}}^{H}}\overbrace{\left\lVert\left[{\bm{P}}^{\top}(\operatorname{diag}(\overline{{\bm{A}}})\otimes\overline{{\bm{B}}}\otimes{\bm{C}}^{\top})\right]_{(H_{\star}+1):H}\right\rVert_{0}}^{\text{eliminating redundant dimensions}}+\overbrace{\underbrace{\left\lVert\left[{\bm{P}}^{\top}\overline{{\bm{A}}}{\bm{P}}\right]_{1:H_{\star},1:H_{\star}}-\overline{{\bm{A}}}_{\star}\right\rVert_{0}}_{\text{aligning the state matrix}}+\underbrace{\left\lVert\left[{\bm{P}}^{\top}(\overline{{\bm{B}}}\otimes{\bm{C}}^{\top})\right]_{1:H_{\star}}-\overline{{\bm{B}}}_{\star}\otimes{\bm{C}}_{\star}^{\top}\right\rVert_{0}}_{\text{aligning input-output interactions}}}^{\text{aligning used dimensions with target model}}.$
@@
Define $\otimes$ as the element-wise product. We can express the target model as:
@@
$=\sum_{m=1}^{n}\operatorname{diag}\left(\begin{bmatrix}\overline{{\bm{A}}}_{\star}&{\bm{O}}\\ {\bm{O}}&{\bm{O}}\\ \end{bmatrix}\right)^{n-m}\left(\begin{bmatrix}{\bm{C}}_{\star}^{\top}\\ {\bm{0}}\end{bmatrix}\otimes\begin{bmatrix}\overline{{\bm{B}}}_{\star}\\ {\bm{0}}\end{bmatrix}\right)x_{m}$ (43)

### v3

@@ C.3 Limitations of Applying Input-injection Methods on SSMs
We start by introducing the necessary notations. Denote the space of S4 mechanisms with $D$ channels as ${\mathcal{F}}_{\text{S4},D}$. Let ${\bm{H}}_{0}=({\bm{h}}_{0}^{(1)},{\bm{h}}_{0}^{(2)},\dots,{\bm{h}}_{0}^{(D)})\in{\mathbb{R}}^{H\times D}$ represent the initial hidden state, and ${\bm{X}}=({\bm{x}}_{1},{\bm{x}}_{2},\ldots,{\bm{x}}_{N})\in{\mathbb{R}}^{D\times N}$ denote the input sequence. The output of the S4 mechanism is represented as $f({\bm{X}};{\bm{H}}_{0})$. Furthermore, for $d$-th channel, let state transition matrix $\overline{{\bm{A}}}^{(d)}=\operatorname{diag}{(a_{1}^{(d)},\cdots,a_{H}^{(d)})}$ and input transition vector $\overline{{\bm{B}}}^{(d)}=(b_{1},\cdots,b_{H})^{\top}$, where $d=1,\ldots,D$. For any vector ${\bm{v}}\in{\mathbb{R}}^{n}$, we use ${\bm{v}}_{i:j}\in{\mathbb{R}}^{j-i}$ to denote the subvector of ${\bm{v}}$ containing elements from $i\in{\mathbb{N}}^{+}$ to $j\in{\mathbb{N}}^{+}$, where $i<j$. Similarly, for any matrix ${\bm{M}}\in{\mathbb{R}}^{m\times n}$, we use ${\bm{M}}_{i_{1}:j_{1},i_{2}:j_{2}}$ to denote the submatrix containing rows $i_{1}\in{\mathbb{N}}^{+}$ to $j_{1}\in{\mathbb{N}}^{+}$ and columns $i_{2}\in{\mathbb{N}}^{+}$ to $j_{2}\in{\mathbb{N}}^{+}$, where $i_{1}<j_{1}$, $i_{2}<j_{2}$.
Proposition 1 (Expressivity of Prefix-Tuning on SSMs).
@@ C.4 Optimal Application of LoRA in SSM-based Models
Several studies (Hu et al., 2023; He et al., 2021) present findings on Transformers, indicating that applying LoRA to linear projection matrices yields performance comparable to or marginally superior to that of attention layers. In contrast, our experimental results on SSMs reveal that applying LoRA to linear projection matrices is more effective than applying it to S6. To elucidate this phenomenon, we examine the influence of updating linear projection matrices on the model’s output.
@@
To make the analysis tractable, we consider a simplified SSM-based architecture composed of the following components:
  • Two input projection matrices: ${\bm{W}}_{\text{in,1}},{\bm{W}}_{\text{in,2}}\in{\mathbb{R}}^{D\times D}$;
@@
  • The S6 module parameterized by diagonal state transition matrices $\{{\bm{A}}^{(d)}\}_{d=1}^{D}$ with ${\bm{A}}^{(d)}\in{\mathbb{R}}^{H\times H}$, the weight matrices $\bm{W}_{{\bm{B}}},\bm{W}_{{\bm{C}}}\in{\mathbb{R}}^{H\times D}$ for computing input-dependent input transition vectors ${\bm{B}}_{n}\in{\mathbb{R}}^{H}$ and output mapping vectors ${\bm{C}}_{n}\in{\mathbb{R}}^{H}$, the down and up projection matrices ${\bm{W}}_{\bm{\Delta},\downarrow}\in{\mathbb{R}}^{D\times R},{\bm{W}}_{\bm{\Delta},\uparrow}\in{\mathbb{R}}^{R\times D}$ (where $R$ is the rank) for low-rank weight matrices for computing the input-dependent step size $\bm{\Delta}_{n}=(\Delta^{(1)}_{n},\ldots,\Delta^{(D)}_{n})\in{\mathbb{R}}_{D}$, for $n=1,\ldots,N$.
Define $\bm{W}_{\text{S6}}=[\bm{W}_{{\bm{B}}}^{\top},\bm{W}_{{\bm{C}}}^{\top},{\bm{W}}_{\bm{\Delta},\uparrow}^{\top}]^{\top}\in{\mathbb{R}}^{(2H+R)\times D}$. In the Mamba implementation, $\bm{W}_{\text{S6}}$ is implemented as the weight matrix of a single linear layer, referred to as x_proj in the codebase. Let the input sequence be ${\bm{X}}=({\bm{x}}_{1},\ldots,{\bm{x}}_{N})\in{\mathbb{R}}^{D\times N}$. At each time step $n$, the S6 module uses two differently projected versions of the input: (i) the input projected via ${\bm{W}}_{\text{in,1}}$ is used to compute the input-dependent parameters $\overline{{\bm{A}}}_{n}$, $\overline{{\bm{B}}}_{n}$, and ${\bm{C}}_{n}$, and (ii) the input projected via ${\bm{W}}_{\text{in,2}}$ serves as the actual input to the S6 module. We note that this formulation generalizes the standard case, which uses a single input projection matrix before the S6 module. In particular, it reduces to the standard case when ${\bm{W}}_{\text{in,1}}={\bm{W}}_{\text{in,2}}$. Then, the output at time step $N$ is given by:
$y^{(d)}_{N}=\underbrace{\overbrace{{\bm{C}}({\bm{W}}_{\text{in,1}}{\bm{x}}_{N})^{\top}}^{\text{Input-dependent }{\bm{C}}_{N}}\sum_{n=1}^{N}\left(\prod_{m=1}^{n}\overbrace{\overline{{\bm{A}}}({\bm{W}}_{\text{in,1}}{\bm{x}}_{m})}^{\text{Input-dependent }\overline{{\bm{A}}}_{m}}\right)\overbrace{\overline{{\bm{B}}}_{n}({\bm{W}}_{\text{in,1}}{\bm{x}}_{n})}^{\text{Input-dependent }\overline{{\bm{B}}}_{n}}}_{\text{Parameters depending on input after projection }{\bm{W}}_{\text{in,1}}}\underbrace{({\bm{W}}_{\text{in,2}}{\bm{x}}_{n})^{(d)}}_{\text{Input after projection}{\bm{W}}_{\text{in,2}}}.$ (24)
To be more specific, the definitions for the relevant terms are:
$\bm{\Delta}_{n}=\operatorname{softplus}({\bm{W}}_{\bm{\Delta},\downarrow}{\bm{W}}_{\bm{\Delta},\uparrow}{\bm{W}}_{\text{in, 1}}\bm{x}_{n}+\bm{\beta}_{\bm{\Delta}}),$ (25)
$\overline{{\bm{A}}}_{n}^{(d)}=\exp(\Delta_{n}^{(d)}{\bm{A}}^{(d)}),\quad\overline{{\bm{B}}}_{n}^{(d)}=\Delta_{n}^{(d)}\bm{W}_{{\bm{B}}}{\bm{W}}_{\text{in, 1}}\bm{x}_{n},\quad{\bm{C}}_{n}=\bm{W}_{{\bm{C}}}{\bm{W}}_{\text{in, 1}}\bm{x}_{n}.$ (26)
When $\bm{\beta}_{\bm{\Delta}}={\bm{0}}$, the output at time step $N$ can be further written as
$y^{(d)}_{N}=(\bm{W}_{{\bm{C}}}{\bm{W}}_{\text{in, 1}}\bm{x}_{n})^{\top}\sum_{n=1}^{N}\left(\prod_{m=1}^{n}\exp(\Delta_{n}^{(d)}{\bm{A}}^{(d)})\right)\Delta_{n}^{(d)}\bm{W}_{{\bm{B}}}{\bm{W}}_{\text{in, 1}}\bm{x}_{n}({\bm{W}}_{\text{in,2}}^{\star}{\bm{x}}_{n})^{(d)},$ (27)
where $\text{where }\bm{\Delta}_{n}=\operatorname{softplus}({\bm{W}}_{\bm{\Delta},\downarrow}{\bm{W}}_{\bm{\Delta},\uparrow}{\bm{W}}_{\text{in, 1}}\bm{x}_{n}).$ (28)
@@
Assume none of the parameters are zero and $D>2H+R$, where $R$ is the rank of ${\bm{W}}_{\bm{\Delta},\downarrow}{\bm{W}}_{\bm{\Delta},\uparrow}$. Lemma 1 in the main body shows that applying LoRA solely to ${\bm{W}}_{\text{in,1}}$ is equivalent to applying it to $\bm{W}_{\text{S6}}$. For completeness, we provide the proof below and restate the lemma for the reader’s convenience.
Lemma 0 (Expressivity of Fine-Tuning Projection Matrices).
Consider two models with the architecture described above. Let:
  • A target model $f^{\star}$ parameterized by ($(\{{\bm{A}}^{\star(d)}\}_{d=1}^{D}$, $\bm{W}_{{\bm{B}}}^{\star}$, $\bm{W}_{{\bm{C}}}^{\star}$, ${\bm{W}}_{\bm{\Delta},\uparrow}^{\star}$, ${\bm{W}}_{\bm{\Delta},\downarrow}^{\star}$, ${\bm{W}}_{\text{in,1}}^{\star}$, ${\bm{W}}_{\text{in,2}}^{\star})$;
  • A frozen model $f_{0}$ parameterized by ($(\{{\bm{A}}^{\star(d)}\}_{d=1}^{D}$, $\bm{W}_{{\bm{B}}}$, $\bm{W}_{{\bm{C}}}$, ${\bm{W}}_{\bm{\Delta},\uparrow}$, ${\bm{W}}_{\bm{\Delta},\downarrow}^{\star}$, ${\bm{W}}_{\text{in,1}}$, ${\bm{W}}_{\text{in,2}}^{\star})$.
The two models share $\{{\bm{A}}^{\star(d)}\}_{d=1}^{D}$, ${\bm{W}}_{\bm{\Delta},\downarrow}^{\star}$, and ${\bm{W}}_{\text{in,2}}^{\star}$, while differing in $\bm{W}_{{\bm{B}}}$, $\bm{W}_{{\bm{C}}}$, ${\bm{W}}_{\bm{\Delta},\uparrow}$, and ${\bm{W}}_{\text{in,1}}$. Then, there exists a projection matrix $\widehat{{\bm{W}}}_{\text{in,1}}$ such that the frozen model matches the output of the target model for any input sequence, i.e.,
$f(\cdot;\{{\bm{A}}^{\star(d)}\}_{d=1}^{D},\bm{W}_{{\bm{B}}},\bm{W}_{{\bm{C}}},{\bm{W}}_{\bm{\Delta},\uparrow},{\bm{W}}_{\bm{\Delta},\downarrow}^{\star},\widehat{{\bm{W}}_{\text{in,1}}},{\bm{W}}_{\text{in,2}}^{\star})=f^{\star}(\cdot;\{{\bm{A}}^{\star(d)}\}_{d=1}^{D},\bm{W}_{{\bm{B}}}^{\star},\bm{W}_{{\bm{C}}}^{\star},{\bm{W}}_{\bm{\Delta},\uparrow}^{\star},{\bm{W}}_{\bm{\Delta},\downarrow}^{\star},{\bm{W}}_{\text{in,1}}^{\star},{\bm{W}}_{\text{in,2}}^{\star}).$ (29)
@@
To prove (29), we substitute (27) into (29), simplify the expression, and show that the equality holds under the following conditions:
$\bm{W}_{{\bm{C}}}^{\star}{\bm{W}}_{\text{in,1}}^{\star}=\bm{W}_{{\bm{C}}}{\bm{W}}_{\text{in,1}}$ (30)
${\bm{W}}_{\bm{\Delta},\uparrow}^{\star}{\bm{W}}_{\text{in,1}}^{\star}={\bm{W}}_{\bm{\Delta},\uparrow}{\bm{W}}_{\text{in,1}}$ (31)
$\bm{W}_{{\bm{B}}}^{\star}{\bm{W}}_{\text{in,1}}^{\star}=\bm{W}_{{\bm{B}}}{\bm{W}}_{\text{in,1}}.$ (32)
Since $\bm{W}_{\text{S6}}=\begin{bmatrix}\bm{W}_{{\bm{B}}}\\ \bm{W}_{{\bm{C}}}\\ {\bm{W}}_{\bm{\Delta},\uparrow}\end{bmatrix}$, the three conditions (30) can be compactly written as
$\bm{W}_{\text{S6}}^{\star}{\bm{W}}_{\text{in,1}}^{\star}=\bm{W}_{\text{S6}}{\bm{W}}_{\text{in,1}}.$ (33)
We now show that for any $\bm{W}_{\text{S6}}$, there exists a matrix ${\bm{W}}_{\text{in,1}}$ that satisfies (33). By applying Singular Value Decomposition (SVD) to $\bm{W}_{\text{S6}}$ , we obtain:
@@
where ${\bm{U}}\in{\mathbb{R}}^{(2H+R)\times(2H+R)}$, ${\bm{\Sigma}}\in{\mathbb{R}}^{(2H+R)\times(2H+R)}$, and ${\bm{V}}\in{\mathbb{R}}^{D\times D}$. The diagonal elements of ${\bm{\Sigma}}$ are in decreasing order. We let
${\bm{W}}_{\text{in,1}}={\bm{V}}\begin{bmatrix}{\bm{\Sigma}}^{-1}{\bm{U}}^{\top}\bm{W}_{\text{S6}}^{\star}{\bm{W}}_{\text{in,1}}^{\star}\\ {\bm{Q}}\end{bmatrix},$ (35)
where ${\bm{Q}}\in{\mathbb{R}}^{(D-2H-R)\times D}$ is an arbitrary matrix to be determined later. Plugging (34) and (35) back to $\bm{W}_{\text{S6}}{\bm{W}}_{\text{in,1}}$ and simplifying results in
$\bm{W}_{\text{S6}}{\bm{W}}_{\text{in,1}}$ (36)
$={\bm{U}}\begin{bmatrix}{\bm{\Sigma}}&{\bm{O}}_{(2H+R)\times(D-2H-R)}\end{bmatrix}{\bm{V}}^{\top}{\bm{V}}\begin{bmatrix}{\bm{\Sigma}}^{-1}{\bm{U}}^{\top}\bm{W}_{\text{S6}}^{\star}{\bm{W}}_{\text{in,1}}^{\star}\\ {\bm{Q}}\end{bmatrix}$ ((34) & (35)) (37)
$=\bm{W}_{\text{S6}}^{\star}{\bm{W}}_{\text{in,1}}^{\star},$ (Simplifying) (38)
@@
Empirical Validation.
To experimentally verify Lemma 1, we conduct a small-scale experiment. Specifically, we train Mamba 130M on three GLUE tasks—RTE, MRPC, and CoLA—for ten epoch under two settings: (1) training only the linear projection layer ($\bm{W}_{\text{in}}$), and (2) training the S6 modules ($\bm{W}_{{\bm{B}}}$, $\bm{W}_{{\bm{C}}}$, ${\bm{W}}_{\bm{\Delta},\uparrow}$). We experiment with various learning rates and, for each configuration, repeated the best-performing setting five times to ensure robustness. As shown in Fig. 3 (training loss) and Table 15 (validation metrics), our results confirm that optimizing only the linear projection layer is as expressive as training the S6 layers. In fact, in all cases, training only the linear projection not only matches, but even outperforms S6 layer training and converges more quickly.
Figure 3: Training cross-entropy loss, with the shaded area indicating the standard deviation, for fine-tuning linear projection layers versus S6 layers. The results show that $\bm{W}_{\text{in}}$ matches the expressivity of $\bm{W}_{{\bm{B}}}$, $\bm{W}_{{\bm{C}}}$, and ${\bm{W}}_{\bm{\Delta},\uparrow}$, while also achieving faster convergence.
Layers RTE MRPC CoLA
$\bm{W}_{\text{in}}$ 69.9 ± 1.2 82.4 ± 0.7 61.1 ± 2.3
$\bm{W}_{{\bm{B}}}$, $\bm{W}_{{\bm{C}}}$, ${\bm{W}}_{\bm{\Delta},\uparrow}$ 68.3 ± 0.8 79.9 ± 1.2 54.9 ± 1.5
Table 15: Mean and standard deviation of validation metrics for fine-tuning linear projection layers versus S6 layers. The results demonstrate that $\bm{W}_{\text{in}}$ effectively captures the expressivity of $\bm{W}_{{\bm{B}}}$, $\bm{W}_{{\bm{C}}}$, and ${\bm{W}}_{\bm{\Delta},\uparrow}$, and achieves superior performance on validation metrics—accuracy for RTE and MRPC, and Matthews correlation coefficient for CoLA.
Appendix D Details of Sec. 5: SDT
@@
Let ${\mathcal{P}}^{H}$ denote the set of all $H\times H$ permutation matrices. Given this formulation, we present our first analysis of parameter efficiency for the S4 model in the following lemma. This analysis is based on the parameters after necessary discretization $(\overline{{\bm{A}}},\overline{{\bm{B}}},{\bm{C}})$. For the reader’s convenience, we restate Lemma 2 below with minor notational changes to facilitate the proof.
Lemma 0 (Minimal Parameter Adjustment for S4 Fine-Tuning).
Consider the parameters after discretization, i.e., $\overline{{\bm{A}}},\overline{{\bm{B}}},{\bm{C}}$. To achieve functional equivalence between the updated model and the target model, i.e., $\hat{f}\equiv f^{\star}$, the minimum number of tunable parameters is:
$\min_{{\bm{P}}\in{\mathcal{P}}^{H}}\overbrace{\left\lVert\left[{\bm{P}}^{\top}(\operatorname{diag}(\overline{{\bm{A}}})\odot\overline{{\bm{B}}}\odot{\bm{C}}^{\top})\right]_{(H_{\star}+1):H}\right\rVert_{0}}^{\text{eliminating redundant dimensions}}+\overbrace{\underbrace{\left\lVert\left[{\bm{P}}^{\top}\overline{{\bm{A}}}{\bm{P}}\right]_{1:H_{\star},1:H_{\star}}-\overline{{\bm{A}}}_{\star}\right\rVert_{0}}_{\text{aligning the state matrix}}+\underbrace{\left\lVert\left[{\bm{P}}^{\top}(\overline{{\bm{B}}}\odot{\bm{C}}^{\top})\right]_{1:H_{\star}}-\overline{{\bm{B}}}_{\star}\odot{\bm{C}}_{\star}^{\top}\right\rVert_{0}}_{\text{aligning input-output interactions}}}^{\text{aligningused dimensions with target model}}.$
@@
Define $\odot$ as the element-wise product. We can express the target model as:
@@
$=\sum_{m=1}^{n}\operatorname{diag}\left(\begin{bmatrix}\overline{{\bm{A}}}_{\star}&{\bm{O}}\\ {\bm{O}}&{\bm{O}}\\ \end{bmatrix}\right)^{n-m}\left(\begin{bmatrix}{\bm{C}}_{\star}^{\top}\\ {\bm{0}}\end{bmatrix}\odot\begin{bmatrix}\overline{{\bm{B}}}_{\star}\\ {\bm{0}}\end{bmatrix}\right)x_{m}.$ (44)